OpenStudy (anonymous):

The reciprocal of two more than a number is three times the reciprocal of the number. Find the number.

OpenStudy (amistre64):

before good notation was developed, they used to define things wordy like this ....

OpenStudy (amistre64):

\[n+2\text{ flips to }\frac{1}{n+2}\] and this is sposed to equate to 3/n \[\frac{1}{n+2}=\frac{3}{n}\]